Please help me solve this problem to the simplest form:
6a^2-21a+15

Answer by atif.muhammad(135)

6a^2 - 21a + 15 = 0
2a^2 - 7a + 5 = 0
2a^2 - 2a - 5a +5 = 0
2a(a-1) -5(a-1) = 0
(2a-5)(a-1)=0
a= 1, 2.5